OpenStudy (pphalke):

two planes leave an airport at the same time, one flying east, the other flying west. the eastbound plane travels 50mph slower. they are 2550 mi apart after 3 hr. find the speed of each plane

OpenStudy (anonymous):

So the westbound planes distance from the start can be described by the following where v is the unknown speed, t is time and d is distance d_west=v*t for for the eastbound plane the speed is v-50, so its distance function will look like the following d_east=(v-50)*t so the total distance between them is the sum of the two d_west+d_west=v*t+(v-50)*t now we can plug in that we know the total distance and time v*3+(v-50)*3=2550 now we can solve this equation v=450 so the westbound plane has a speed of 450 and the eastbound a speed 50 mph lower that is 400

OpenStudy (damonte212):

..,,.